OpenStudy (kitkat16):

help please! will give medal! The equation of a circle is x2 + y2 + Cx + Dy + E = 0. If the radius of the circle is decreased without changing the coordinates of the center point, how are the coefficients C, D, and E affected?

OpenStudy (anonymous):

Convert the above equation to form\[(x-\alpha)^2+(y-\beta)^2=r^2\] from there it is easy to note that \[C=-\alpha*2, D=-\beta*2\] and \[\alpha^2+\beta^2-r^2 = E\]; where r represents the radius of the circle and \[\alpha, \beta\] are the coordinate of the center of the circle.

OpenStudy (anonymous):

Do not use sentences like giving medal while asking questions for the love of god. Its a forum and not Award distribution ceremony.

OpenStudy (unklerhaukus):

complete the squares

OpenStudy (kitkat16):

Sorry about the medal comment. I'm new and this is the third question I have ask and this is the first time anyone responded. I saw other people writing that. I didn't know. E =R^2 and it decreases so it changes. The center point doesn't change so C & D don't change C and D are unchanged but E increases

OpenStudy (anonymous):

Yes, you are correct

OpenStudy (kitkat16):

thanks for responding :) I don't even know why they put me in this class. I've never done this stuff. Thanks again!

OpenStudy (anonymous):

Which class are you referring to?

OpenStudy (anonymous):

E increases? Wtf? No.. thats NOT even one of the answers to chose from.. the fnck outta here with them false answers.. Tired of failing cause people put up wrong answers. -_______-

OpenStudy (anonymous):

it's no ones fault but yours ^

OpenStudy (radmeout):

lmao^